Shiftsum for EVE Jan 8 2000


Subject: Shiftsum for EVE Jan 8 2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:11:07 CST


                Evening Shift 8 Jan 00
 
On Shift: SL,RF

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E799 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0.1
Accounting of lost time:
  KFRENDs, Pipeline reset, starting and stopping runs,change tapes

========================================================================
Tapes brought to FCC this shift: UP* 107-109
                                 UPS 103
========================================================================
Accesses during shift: None

========================================================================
Runs ended this shift:

15441 E799 data 21M events
15442 Muon run, 100K events

 
=======================================================================
Overview of shift: Fairly smooth E799 data. Crate 2 reboot, Several trigger
hang-ups, resolved with multiple KFRENDs. Pipeline messages starting
around 10pm.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15441 Run ongoing as shift starts.

        240 1538 Trigger hangs. Many kfrends and cycle repeater
                        doesn't help. Rest pipeline, crate 2 doesn't
                        init. finally hard reboot of crate 2 solves problem.

                1624 Trig. hangs again; multiple kfrends does trick.

                1932 Tapes full. end run 15441

15442 1952 Start muon run; change NM2SR pol. to pos.
                        L3 accept fraction warning; Outbox problems
                        Muon run ends. DRT_FRESH_START.

15443 2025 Start BEAM_799 run with NM4AN NEG.

                2200 "Message from pipeline {}"

                2215 "Message from pipeline {}" again Wide pedestal
                        error. BobT paged. Says its a monitoring problem.
                        We should try to reset pipeline when beam is down
                        for 15 min. or more. It may need a crate reboot.
                
========================================================================
Items to Follow up on Next Shift
 
  Keep an eye on DCOR LV power supply for DC2, it is reading low (4.8V).
  
  Label 30 QKE tapes for short reg run and other special runs
 
  Don't forget to hit "reset quartz" button after each KFREND.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:11:09 CST